OpenStudy (anonymous):

A man of weight W is standing on a lift which is moving upward with an acceleration a. The apparent wieght of the man is

OpenStudy (anonymous):

See, I think the new weight would be: W1 = m(g + a) = mg + ma = W + ma = W + (W/g)a Please confirm with other experts also.

OpenStudy (anonymous):

Yes your answer is right

OpenStudy (anonymous):

Can you explain it in much more detail?

OpenStudy (anonymous):

Thanks Yaar.

OpenStudy (anonymous):

How you got m(g+a)?

OpenStudy (anonymous):

see, when the object is moving upward, always add the acceleration due to gravity and the new acceleration. Multiply the sum by mass of the object. U will get the force acting on it.

OpenStudy (anonymous):

The force acting on it would be the aapparent wieght of man?

OpenStudy (anonymous):

apparent

OpenStudy (anonymous):

Will it be the apparent weight of man?

OpenStudy (anonymous):

I think most probably it should be.

OpenStudy (anonymous):

Ok, Thanks a lot

OpenStudy (anonymous):

U r welcome.

OpenStudy (anonymous):

yes u have given the right formula but how? here it is when the lift is moving up, the normal force is graeter than the weight of man mg but apparent weight or weight in general of a body is determined by the NORMAL FORCE/NORMAL REACTION hence here N-mg=ma N=mg+ma again i repeat when we say the WEIGHT OF A BODY IS SOMETHING IT IS ACTUALLY THE NORMAL REACTION/FORCE AND NOT mg of the body